One railroad corridor with the following traffic and operating characteristics:

80-Car Trains

16 Trains per Day

Cars on each train are 286,000 lb GRL railcars (50 cars are fully loaded, 30 are fully empty)

The railcar Light Weight = 34,500 lbs

Number of Locomotives per Train = 2

Locomotive Weight = 425,000 lbs. each

Length of Route = 1,000 miles

One year has 365 days

1 ton = 2,000 lbs

GRL = maximum total weight when a railcar is fully loaded (maximum lading plus light weight of car)

1) What is the annual gross tonnage (in millions) that any one point on the line would experience?

2) What is the total annual revenue tonnage (in millions) that any one point on the line would experience?

3) How many billion gross ton-miles does the corridor accrue annually?

4) How many billion revenue ton-miles does the corridor accrue annually?

5) What are the total annual train-miles for the corridor (in millions)?
Answer #1

1) Annual gross tonnage (in millions):-

Gross tonnage(including locomotive) = weight of lading + rail car(light weight) + locomotive

GRL/car =286,000 lb

so, Lading/car = 286,000 - 34500 = 251500 lbs

Now, Gross tonnage/train = 50 x GRL + 30 x light weight + 2 x Locomotive weight

                                           = 50 x 286000 + 30 x 34500 + 2 x 425000 = 16185000 lbs

16 trains/day = 16185000 x 16 = 258960000 lbs

Annual gross tonnage(in millions) = (258960000 x 365)/(2000 x 106) =      47.2602 Million Gross Tons.

Answer = 47.2602 MGT

2) Annual revenue tonnage (in millions) :-

Revenue tonnage = weight of lading

Annual revenue tonnage (in millions) = (16 x (50x 251500) x 365) / (2000 x 106) = 36.719 Million revenue tons

3) Billion gross ton miles annually -

Annual Gross ton miles(including locomotive) = (16 x (50 x 286000 + 30 x 34500 + 2 x 425000) x 1000 miles) x 365 / (2000 x 109) = 47.2602 Billion Gross tons

5) Annual train -miles for corridor ( in millions )

= 365 x 16 x 1000 x (1/106 ) = 5.84 million train- miles.